OpenStudy (anonymous):

The function f(x) is shown below. f(x) = (3)x - 2 The function f(x) is shifted to the left by 13 units. Which of the following best represents the new function? f(x) = (3)x + 11 f(x) = (3)x + 13 f(x) = (3)x -15 f(x) = (3)x - 13

OpenStudy (anonymous):

pgpilot so would it be fx = (3) x + 13 or + 11 becuase of the - 2

OpenStudy (anonymous):

you have to account for the -2

OpenStudy (anonymous):

so it would be + 11 becuase of the -2 right?

OpenStudy (akashdeepdeb):

TO SHIFT TO THE RIGHT, you subtract the number of units to (x-a) as f(x-a). TO SHIFT IT TO LEFT, you add the number of units to (x+a) as f(x+a). TO SHIFT IT UP, you add the number of units to f(x). TO SHIFT DOWN, you subtract the number of units to f(x).

OpenStudy (anonymous):

f(x)=3(x+13)-2

OpenStudy (anonymous):

pgpilot am i correct with the x + 11 becuase of the - 2

OpenStudy (anonymous):

it looks like your function is \(f(x)=3x-2\), correct? if so, to find the shift left of 13 units you would need to compute \(g(x)=f(x+13)=3(x+13)-2 = 3x+39-2=3x+37\) but i don't see that option in your choices. to prove to you that this is what it should be let's look at the x-intercept. for \(f(x)\) the x-intercept is 2/3. Now move this to the left by 13 and you'll get \[x=-12\frac{1}{3}=-\frac{37}{3}\] and what's the x-intercept of g(x)? well, it's -37/3... just as it should be.

OpenStudy (anonymous):

maybe this might clarify things for you The function f(x) is shown below. f(x) = (3)^x - 2 The function f(x) is shifted to the left by 13 units. Which of the following best represents the new function? f(x) = (3)^x + 11 f(x) = (3)^x + 13 f(x) = (3)^x -15 f(x) = (3^)x - 13

OpenStudy (anonymous):

you would still do the same thing... \[g(x)=f(x+13) =3^{x+13}-2\]

OpenStudy (anonymous):

you sure int's not \[f(x)=3^{x-2}\]

OpenStudy (anonymous):

that is what it is

OpenStudy (anonymous):

thats why i put the ^ in front of x-2

OpenStudy (anonymous):

yes, but you needed to put parentheses around the x-2 to be clear. 3^x-2 = \(3^x-2\) whereas 3^(x-2) = \(3^{x-2}\) so you still do the same thing but now the -2 is in the exponent as well, so you get \[g(x)=f(x+13)=3^{(x+13)-2} = 3^{x+13-2}=3^{x+11}\]

OpenStudy (anonymous):

ok thank you and sorry can you help me with another question

OpenStudy (anonymous):

sure just @me in the post.

OpenStudy (anonymous):

Evaluate cube root of 5 multiplied by square root of 5 over cube root of 5 to the power of 5

OpenStudy (anonymous):

@pgpilot326 Evaluate cube root of 5 multiplied by square root of 5 over cube root of 5 to the power of 5

OpenStudy (anonymous):

please put in a new post. thanks